OpenStudy (anonymous):

A pyramid has a volume of 16 and an altitude of 5. What is the area of it's base?

OpenStudy (anonymous):

V = (1/3)*A*h 16 = (1/3)*A*5 16*3 = 5A 48 = 5A 48/5 = A 9.6 = A A = 9.6 So the area of the base is 9.6 square units.